March 6, 1996

Miss Iusta Mo 34 Winchester Road Sheffield England

Dear Miss Mo,

This letter is to inform you that you obtained grade A in your recent examinations.

Mr Arthur Minit 43 Sheffield Road Winchester England

Dear Mr Minit,

This letter is to inform you that you obtained grade C in your recent examinations.